# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

### v1

@@ 3.3. The moving GCD theorem for parabolic Riemann surfaces
One of the key components in [9] is the extension of the GCD theorem in [14] to the case where the coefficients of two coprime polynomials are allowed to be meromorphic functions of the same growth. In this subsection, we aim to extend this line of argument to the setting of parabolic Riemann surfaces. Before we proceed, let us briefly outline the development of the GCD theorem.
@@
To proceed, we will first derive a version of the second main theorem with moving targets of the same growth, which will serve our purpose.
Let $\mathcal{Y}$ be a parabolic open Riemann surface. Let $a_{0},\ldots,a_{n}$ be meromorphic functions, not all zeros, on $\mathcal{Y}$, and let $L:=a_{0}X_{0}+\dots+a_{n}X_{n}$. Then $L$ defines a hyperplane $H$ in $\mathbb{P}^{n}$, over the field generated by $a_{0},\ldots,a_{n}$ over $\mathbb{C}$. We note that $H(z)$ is the hyperplane determined by the linear form $L(z)=a_{0}(z)X_{0}+\dots+a_{n}(z)X_{n}$ for $z\in\mathcal{Y}$ where $a_{0},\cdots,a_{n}$ are entire and there is no common zero of $a_{0},\cdots,a_{n}$ at $z$. The definition of the Weil function, proximity function and counting function can be easily extended to moving hyperplanes. For example,
@@
for $z\in\mathcal{Y}$ which is not a common zero of $a_{0},\cdots,a_{n}$, or a pole of any $a_{k}$, $0\leq k\leq n$. The First Main Theorem for a moving hyperplane $H$ can be stated as
(3.4) $T_{\mathbf{f},\mathcal{Y}}(r)=N_{\mathbf{f},\mathcal{Y}}(H,r)+m_{\mathbf{f},\mathcal{Y}}(H,r)+T_{\mathbf{a},\mathcal{Y}}(r)+O(1),$
where ${\mathbf{a}}:=[a_{0}:\cdots:a_{n}]$.
Without loss of generality, we will normalize the linear forms $L_{j}$, $1\leq j\leq q$, such that for each $1\leq j\leq q$, there exists $0\leq j^{\prime}\leq n$ with $a_{ij^{\prime}}=1$. Let $t$ be a positive integer and let $V(t)$ be the complex vector space spanned by the elements
@@
Let $V$ be a $\mathbb{C}$-vector space spanned by finitely many meromorphic functions. We say that a holomorphic map ${\mathbf{f}}=(f_{0},\dots,f_{n}):\mathcal{Y}\to\mathbb{P}^{n}({\mathbb{C}})$ is linearly non-degenerate over $V$ if whenever we have a linear combination $\sum_{i=1}^{m}a_{i}f_{i}=0$ with $a_{i}\in V$, then $a_{i}=0$ for each $i$. Otherwise we say that ${\mathbf{f}}$ is linearly degenerate over $V$.
@@
[…]
(3.6) $\int_{S(r)}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}(z)}({\mathbf{f}}(z))d\mu_{r}+\frac{1}{u}N_{W}(0,r)$
(3.7) $\leq_{\operatorname{exc}}\left(\frac{w}{u}(n+1)+\epsilon\right)T_{\mathbf{f},\mathcal{Y}}(r)+\frac{w}{u}(n+1)(t+2)\max_{1\leq j\leq q}T_{L_{j},{\mathcal{Y}}}(r)$
(3.8) $+{1\over 2}(n+1)^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
@@
where $T_{L_{j},{\mathcal{Y}}}(r)=T_{[a_{j0}:\cdots:a_{jn}],{\mathcal{Y}}}(r)$, $w=\dim V(t+1)$, $u=\dim V(t)$, and the maximum is taken over all subsets $J$ of $\{1,\dots,q\}$ such that $H_{j}(z)$, $j\in J$, are in general position.
Using the above theorem, and through the same argument from the proofs of [14, Theorem 5.7], [8, Theorem 26], and [9, Theorem 4.4], we can get the following GCD theorem (note we now consider the case that $u_{1},\ldots,u_{n}$ are units on $\mathcal{Y}$). We first fix some notations: Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$. Assume that one of the coefficients in each expansion of $F$ and $G$ is 1. For every positive integer $t$, we denote by $V_{F,G}(t)$ the (finite-dimensional) $\mathbb{C}$-vector space spanned by $\prod_{\alpha}\alpha^{n_{\alpha}}$, where $\alpha$ runs through all non-zero coefficients of $F$ and $G$, $n_{\alpha}\geq 0$ and $\sum n_{\alpha}=t$; we also put $d_{t}:=\dim V_{F,G}(t)$. For every integer $m\geq d$, we let $M=M_{m,n,d}=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$. We now define the characteristic function of a polynomial over $K$. Let $Q=\sum_{{\mathbf{i}}\in I_{Q}}a_{\mathbf{i}}\cdot{\mathbf{x}}^{\mathbf{i}}\in K[x_{1},\ldots,x_{n}]$, where ${\mathbf{i}}=(i_{1},\ldots,i_{n})$, ${\mathbf{x}}^{\mathbf{i}}=x_{1}^{i_{1}}\ldots x_{n}^{i_{n}}$, and $a_{\mathbf{i}}\neq 0$ if ${\mathbf{i}}\in I_{Q}$. We let
(3.9) $\|Q\|_{z}:=\max_{{\mathbf{i}}\in I_{Q}}\{|a_{\mathbf{i}}(z)|\},$
and define the characteristic function of $Q$ as
(3.10) $T_{Q,\mathcal{Y}}(r):=T_{[\cdots:a_{\mathbf{i}}:\cdots],\mathcal{Y}}(r),$
where $a_{\mathbf{i}}$ is taken for every ${\mathbf{i}}\in I_{Q}$.
Let $n\geq 2$ be a positive integer. Let $u_{1},\ldots,u_{n}$ be units on $\mathcal{Y}$, i.e. holomorphic functions on $\mathcal{Y}$ without zeros and let ${\bf u}=(u_{1},\ldots,u_{n})$. Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$. Assume that one of the coefficients in each expansion of $F$ and $G$ is 1. Then there exists a positive constant $c$ and a positive integer $b$ such that if the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, then, for any $\delta>0$, we have
$N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1+\frac{\epsilon}{M}\right)mn\max_{1\leq i\leq n}T_{u_{i},{\mathcal{Y}}}(r)$
$+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
where $w=d_{Mb}$, $u=d_{M(b-1)}$, $c^{\prime}=\frac{w}{u}M(b+2)+1$, and $c^{\prime\prime}=M\frac{w^{2}}{u}$, and $M^{\prime}=\binom{m+n}{n}-M$.
We outline a proof here.
Proof.
${\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}\in(F,G)_{m}$
for some choice $c_{z,{\bf j}}\in K$. Then, for each such ${\bf i}$, there is a linear form over $K$ such that
We note that since there are correspondingly only finitely many choices of a monomial basis of $V_{m}$, there are only finitely many choices of $c_{z,{\bf j}}$, even as $z$ ranges over all of $\mathbb{C}$. Following the explicit construction in the proof of [9, Theorem 4.3], we can find a nonzero element $c_{z}\in V_{F,G}(M)$ such that the normalized linear form $L_{z,{\bf i}}:=c_{z}\tilde{L}_{z,{\bf i}}$ satisfies
and
(3.11) $T_{L_{z,{\bf i}},{\mathcal{Y}}}(r)\leq M(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)),$
by [9, (4.12)]. Moreover, the explicit construction also demonstrates that there are only a finitely many choice for $c_{z}$, even though $z$ runs through all of $\mathbb{C}$
Following the steps in the proof of Theorem 4.3 in [9], we then obtain
Consequently, we have
(3.13) $\int_{S(r)}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}(\log|L_{z,{\bf i}}(\Phi({\bf u})|-\log\|L_{z,{\bf i}}\|)d\mu_{r}\leq M^{\prime}mnT_{{\bf u},\mathcal{Y}}(r)+M(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+O(1).$
Also, from (3.3) in the proof of [14, Theorem 5.7], we have
(3.14) $\int_{S(r)}\log\|\Phi({\bf u})\|d\mu_{r}=T_{\Phi({\bf u}),{\mathcal{Y}}}(r)+N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)+O(1).$
Since the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, we have that $\Phi({\bf u}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{M-1}$ is linearly nondegenerate over $V_{F,G}(Mb)=V(b)$. We note that, when computing Weil functions, we may assume that the coefficients of $L_{z,{\bf i}}$ are entire functions without common zeros. By this construction, for each $z\in{\mathcal{Y}}$, the hyperplanes defined by evaluating the linear forms $L_{z,{\bf i}}$ at $z$ with $|{\bf i}|=m$ and ${\bf i}\not\in I_{z}$ are in general position. Additionally, we observe that (3.11) and (3.13) remain valid under this choice of linear forms. Applying Theorem 14 to $\Phi({\bf u})$ and the set of the linear forms $\{L_{z,{\bf i}}:|{\bf i}|=m,{\bf i}\not\in I_{z}\}$, we have the following inequality:
(3.15) $\int_{S(r)}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))d\mu_{r}+\frac{1}{u}N_{W}(0,r)$
(3.16) $\leq_{\operatorname{exc}}\left(\frac{w}{u}M+\epsilon\right)T_{\Phi({\bf u}),\mathcal{Y}}(r)+\frac{w}{u}M^{2}(b+2)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))$
(3.17) $+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
where
$\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))=-\log|L_{z,{\bf i}}(\Phi({\bf u})|+\log\|\Phi({\bf u})\|+\log\|L_{z,{\bf i}}\|+O(1).$
By combining (3.15) with (3.13), (3.14), we get
$MN_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)+\frac{1}{u}N_{W}(0,r)$
$\leq_{\operatorname{exc}}\left(\frac{w}{u}M-M+\epsilon\right)T_{\Phi({\bf u}),\mathcal{Y}}(r)+M^{\prime}mnT_{{\bf u},\mathcal{Y}}(r)$
$+\left(\frac{w}{u}M^{2}(b+2)+M\right)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
which proves the theorem. ∎
@@
(3.19) $u:=\dim V(t)\leq\left(\frac{1}{\epsilon}\right)^{qn}.$
Consequently,
(3.20) $w:=\dim V(t+1)\leq(1+\epsilon)\left(\frac{1}{\epsilon}\right)^{qn}.$
Therefore, the constants appeared in (3.6) depend only on $\epsilon$ and the number of hyperplanes.
@@
Let $n\geq 2$ be a positive integer. Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F$ and $G$ be coprime polynomials in $K[x_{1},\ldots,x_{n}]$. Then, for any $\epsilon>0$, there exist a positive integer $m$ and positive reals $c_{i}$, $1\leq i\leq 4$, all depending only on $\epsilon$, $n$, and the degree of $F$ and $G$ such that for all $n$-tuple $(u_{1},\ldots,u_{n})$ of units on $\mathcal{Y}$ satisfying
(3.21) $\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)\geq c_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have either
(3.22) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}c_{2}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
@@
(3.23) $N_{\rm gcd,\mathcal{Y}}(F(u_{1},\ldots,u_{n}),G(u_{1},\ldots,u_{n}),r)\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c_{4}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
@@
We note that $c_{i},1\leq i\leq 4,$ will be determined during the course of the proof. As in the proof of [9, Theorem 4.1], we will assume that one of the coefficients in each expansion of $F$ and $G$ is 1. Recall that $M:=M_{m}:=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$ and $M^{\prime}:=M^{\prime}_{m}:=\binom{m+n}{n}-M=O(m^{n-2})$.
Let $\epsilon>0$. To establish (3.22) or (3.23), we can assume that $\epsilon$ is sufficiently small. Based on the above estimates, we can choose a real $C_{1}\geq 1$ independent of $\epsilon$ and the $u_{i},1\leq i\leq n,$ such that $m=C_{1}\epsilon^{-1}\geq 2d$, and

### v2

@@ 3.4. The moving GCD theorem in the parabolic setting
One of the key components in [9] is the extension of the GCD theorem in [13] to the case where the coefficients of two coprime polynomials are allowed to be meromorphic functions of the same growth. In this subsection, we aim to extend this line of argument to the parabolic setting. Before we proceed, let us briefly outline the development of the GCD theorem.
@@
To proceed, we will first derive a version of the second main theorem with moving targets of the same growth, which will serve our purpose. See [18] or [17, Chapter 6].
Let $a_{0},\ldots,a_{n}$ be meromorphic functions, not all zeros, on $\mathbb{C}$}, and let $L:=a_{0}X_{0}+\dots+a_{n}X_{n}$. Then $L$ defines a hyperplane $H$ in $\mathbb{P}^{n}$, over the field generated by $a_{0},\ldots,a_{n}$ over $\mathbb{C}$. We note that $H(z)$ is the hyperplane determined by the linear form $L(z)$. The definition of the Weil function, proximity function and counting function can be easily extended to moving hyperplanes. For example,
@@
for $z\in\mathcal{Y}$ which is not a common zero of $a_{0},\cdots,a_{n}$, or a pole of any $a_{k}$, $0\leq k\leq n$. We now consider a family of linear forms $L_{j}:=\sum_{i=0}^{n}a_{ij}X_{j}$, $1\leq j\leq q$, as the aformentioned linear forms. Without loss of generality, we normalize the linear forms $L_{j}$, $1\leq j\leq q$, such that for each $1\leq j\leq q$, there exists $0\leq j^{\prime}\leq n$ with $a_{jj^{\prime}}=1$. Let $t$ be a positive integer and let $V(t)$ be the complex vector space spanned by the elements
@@
Let $V$ be a $\mathbb{C}$-vector space spanned by finitely many meromorphic functions. We say that a holomorphic map $f=(f_{0},\dots,f_{n}):{\mathbb{C}}\to\mathbb{P}^{n}({\mathbb{C}})$ is linearly independent over $V$ if whenever we have a linear combination $\sum_{i=0}^{n}a_{i}f_{i}=0$ with $a_{i}\in V$, then $a_{i}=0$ for each $i$. Otherwise we say that $f$ is linearly dependent over $V$.
@@
Let $f=(f_{0},\ldots,f_{n}):{\mathbb{C}}\to\mathbb{P}^{n}(\mathbb{C})$ be a holomorphic map where $f_{0},f_{1},\ldots,f_{n}$ are entire functions without common zeros. Let $H_{j}$, $1\leq j\leq q$, be arbitrary (moving) hyperplanes given by $L_{j}:=a_{j0}X_{0}+\dots+a_{jn}X_{n}$ where $a_{j0},\cdots,a_{jn}$ are meromorphic functions, not all zeros on $\mathbb{C}$}. Let $t$ be a positive integer, $1=b_{1},b_{2},\cdots,b_{u}$ be a basis of $V(t)$ and $b_{1},\cdots,b_{w}$ a basis of $V(t+1)$. Denote by $W$ the Wronskian of $\{hb_{m}f_{k}\,|\,1\leq m\leq w,\,0\leq k\leq n\}$, where $h$ is a meromorphic function such that $hb_{1},\ldots,hb_{w}$ are entire functions without common zeros. Let $\mathcal{Y}\subset{\mathbb{C}}$ be as above. If $f$ is linearly non-degenerate over $V(t+1)$, then, we have the following inequality
(3.22) $\int_{S^{\sigma}_{r}}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}(z)}(f(z))d\mu_{r}+\frac{1}{u}N_{W,\sigma}(0,r)$
(3.23) $\leq_{\operatorname{exc}}\left(\frac{w}{u}(n+1)\right)T_{f,\sigma}(r)+\frac{w}{u}(n+1)(t+2)\sum_{j=1}^{q}T_{L_{j},\sigma}(r)$
(3.24) $+\frac{w}{2u}(n+1)(nw+w-1)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+\frac{1}{u}c_{{1\over W},\sigma}(r)+O(\log T_{f}(r)),$
@@
where $T_{L_{j},\sigma}(r)=\sum_{i=0}^{n}T^{M}_{a_{ji},\sigma}(r)$, $w=\dim V(t+1)$, $u=\dim V(t)$, and the maximum is taken over all subsets $J$ of $\{1,\dots,q\}$ such that $H_{j}(z)$, $j\in J$, are in general position.
@@
(3.26) $u:=\dim V(t)\leq\left(\frac{1}{\epsilon}\right)^{qn},~~~\mbox{and consequently},~~~~w:=\dim V(t+1)\leq(1+\epsilon)\left(\frac{1}{\epsilon}\right)^{qn}.$
Using the above theorem, and through the same argument from the proofs of [13, Theorem 5.7], [8, Theorem 26], and [9, Theorem 4.4], we can get the following GCD theorem (note we now consider the case that $u_{1},\ldots,u_{n}$ are units on $\mathcal{Y}$). We first fix some notation. Let $K$ be a subfield of the field of meromorphic functions on $\mathbb{C}$}. Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$. Assume that one of the coefficients in the expansions of $F$ and $G$ is equal to 1. For every positive integer $t$, we denote by $V_{F,G}(t)$ the (finite-dimensional) $\mathbb{C}$-vector space spanned by $\prod_{\alpha}\alpha^{n_{\alpha}}$, where $\alpha$ runs through all non-zero coefficients of $F$ and $G$, $n_{\alpha}\geq 0$ and $\sum n_{\alpha}=t$; we also put $d_{t}:=\dim V_{F,G}(t)$. For every integer $m\geq d$, we let $M=M_{m,n,d}=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$. Let $Q=\sum_{{\mathbf{i}}\in I_{Q}}a_{\mathbf{i}}\cdot{\mathbf{x}}^{\mathbf{i}}\in K[x_{1},\ldots,x_{n}]$, where ${\mathbf{i}}=(i_{1},\ldots,i_{n})$, ${\mathbf{x}}^{\mathbf{i}}=x_{1}^{i_{1}}\ldots x_{n}^{i_{n}}$, and $a_{\mathbf{i}}\neq 0$ if ${\mathbf{i}}\in I_{Q}$. We define the parabolic characteristic function of $Q$ as
(3.27) $T_{Q,\sigma}(r):=\sum_{{\mathbf{i}}\in I_{Q}}T^{M}_{a_{\mathbf{i}},\sigma}(r).$
Let $n\geq 2$ be a positive integer. Let $K$ be a subfield of the field of meromorphic functions on $\mathbb{C}$}. Let $\mathcal{Y}\subset{\mathbb{C}}$ be as above. Let ${\bf u}=(u_{1},\ldots,u_{n})$, where $u_{1},\ldots,u_{n}$ are meromorphic functions on $\mathbb{C}$} such that each $u_{i}|_{\mathcal{Y}},1\leq i\leq n,$ is a unit (i.e. without zeros and poles on $\mathcal{Y}$}). Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$. Assume that at least one of the coefficients in the expansions of $F$ and $G$ is equal to 1. Let $m\geq d$ and $b$ be a positive integer. If the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, then we have
$N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1\right)mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r)+O(\max_{1\leq i\leq n}\log T_{u_{i}}(r))$
$+c^{\prime}(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r+c_{1/h,\sigma}(r)+c_{1/W,\sigma}(r)),$
where $h$ is meromorphic function such that $F({\bf u})/h$ and $G({\bf u})/h$ are entire and have no common zero, and $W$ is the Wronskian which arises when applying Theorem 18, $w=d_{Mb}$, $u=d_{M(b-1)}$, $c^{\prime}=\frac{w}{u}M(b+3)+2,$ and $c^{\prime\prime}=M{w^{2}\over u}$, and $M^{\prime}=\binom{m+n}{n}-M$.
We outline a proof here.
Proof.
For a subset $T\subset K[\mathbf{x}]$, where $\mathbf{x}:=(x_{1},\ldots,x_{n})$, we denote
$T_{m}=\{f\in T\,:\deg f\leq m\}.$
${\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}\in(F,G)_{m}$
for some choice $c_{z,{\bf j}}\in K$. Then, for each such $\bf i$}, there is a linear form $\tilde{L}_{z,{\bf i}}$ over $K$ such that
We note that since there are correspondingly only finitely many choices of a monomial basis of $V_{m}$, there are only finitely many choices of $c_{z,{\bf j}}$, even as $z$ ranges over all of $\mathcal{Y}$}. Following the explicit construction in the proof of [9, Theorem 4.3], we can find a nonzero element $c_{z}\in V_{F,G}(M)$ such that the normalized linear form $L_{z,{\bf i}}:=c_{z}\tilde{L}_{z,{\bf i}}$ satisfies
and
(3.28) $T_{L_{z,{\bf i}},\sigma}(r)\leq M(T_{F,\sigma}(r)+T_{G,\sigma}(r)),$
by [9, (4.12)]. Moreover, the explicit construction also demonstrates that there are only a finitely many choice for $c_{z}$, even though $z$ runs through all of $\mathcal{Y}$}.
Following the steps in the proof of Theorem 4.3 in [9], we then obtain
Since the functions $u_{i}$ are units on $\mathcal{Y}$, and that
(3.30) $\log^{+}|{\bf u}(z)^{\bf i}|\leq\sum_{j=1}^{n}i_{j}\log^{+}|u_{j}(z)|\leq m\sum_{j=1}^{n}\log^{+}|u_{j}(z)|,$
we have
(3.31) $T^{M}_{{\bf u}^{\bf i},\sigma}(r)\leq m\sum_{j=1}^{n}T^{M}_{u_{j},\sigma}(r)\leq mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r).$
Consequently, by (3.29), we obtain
(3.32) $\int_{S^{\sigma}_{r}}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}(\log|L_{z,{\bf i}}(\Phi({\bf u})|-\log\|L_{z,{\bf i}}\|_{z})d\mu_{r}\leq M^{\prime}mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r)+M(T_{F,\sigma}(r)+T_{G,\sigma}(r))+O(1).$
Note that the map $\phi_{1}({\bf u}),\dots,\phi_{M}({\bf u})$ may not be a reduced presentation of $\Phi({\bf u})$. Let $h$ be a meromorphic function such that $F({\bf u})/h$ and $G({\bf u})/h$ are entire and have no common zero. Let $\psi_{i}:=\phi_{i}({\bf u})/h$ for $i=1,\dots,M$. Then $\Psi:=(\psi_{1},\dots,\psi_{M})$ is a reduced representation of $\Phi({\bf u})$. Then
(3.33) $\int_{S^{\sigma}_{r}}\log\|\Phi({\bf u})\|d\mu_{r}=\int_{S^{\sigma}_{r}}\log\|\Psi\|d\mu_{r}+\int_{S^{\sigma}_{r}}\log|h|d\mu_{r}=T_{\Psi,\sigma}(r)+\int_{S^{\sigma}_{r}}\log|h|d\mu_{r}.$
On the other hand, since $u_{i}$ are units on $\mathcal{Y}$, the poles of $h$ only arises from the poles of coefficients of $F$ and $G$. More precisely,
(3.34) $N_{h,\sigma}(\infty,r)\leq T_{F,\sigma}(r)+T_{G,\sigma}(r)$
by applying the inequality of the first main theorem in Theorem 11. It is also clear that this choice of $h$ gives
(3.35) $N_{h,\sigma}(0,r)=N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r).$
By Lemma 9, (3.34) and (3.35), we have
(3.36) $\int_{S^{\sigma}_{r}}\log|h|d\mu_{r}$ $=N_{h,\sigma}(0,r)-N_{h,\sigma}(\infty,r)+c_{h,\sigma}(r)$
(3.37) $\geq N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)-\left(T_{F,\sigma}(r)+T_{G,\sigma}(r)\right)+c_{h,\sigma}(r).$
Then we can derive from (3.33) that
(3.38) $\int_{S^{\sigma}_{r}}\log\|\Phi({\bf u})\|d\mu_{r}\geq T_{\Psi,\sigma}(r)+N_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)+c_{h,\sigma}(r)-\left(T_{F,\sigma}(r)+T_{G,\sigma}(r)\right).$
Since the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, it follows that $\Phi({\bf u}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{M-1}$ is linearly independent over $V_{F,G}(Mb)=V(b)$. We note that, when computing Weil functions, we have the freedom to assume that the coefficients of $L_{z,{\bf i}}$ are entire functions without common zeros, by multiplying a suitable meromorphic function. By this construction, for each $z\in{\mathcal{Y}}$, the hyperplanes defined by evaluating the linear forms $L_{z,{\bf i}}$ at $z$ with $|{\bf i}|=m$ and ${\bf i}\not\in I_{z}$ are in general position. For the computation (3.28) and (3.32), however, we will use the original $L_{z,{\bf i}}$. By applying Theorem 18 to $\Phi({\bf u})$ and the set of the linear forms $\{L_{z,{\bf i}}:|{\bf i}|=m,{\bf i}\not\in I_{z}\}$, we have the following inequality:
(3.39) $\int_{S^{\sigma}_{r}}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))d\mu_{r}+\frac{1}{u}N_{W,\sigma}(0,r)$
(3.40) $\leq_{\operatorname{exc}}\left(\frac{w}{u}M+\epsilon\right)T_{\Psi,\sigma}(r)+\frac{w}{u}M^{2}(b+2)(T_{F,\sigma}(r)+T_{G,\sigma}(r))$
(3.41) $+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+{1\over u}c_{1/W,\sigma}(r)+O(\log\max_{1\leq j\leq n}T_{u_{j}}(r)),$
where
$\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))=-\log|L_{z,{\bf i}}(\Phi({\bf u}))|+\log\|\Phi({\bf u})\|+\log\|L_{z,{\bf i}}\|+O(1),$
and $W$ is the wronskian of $\psi_{1},\dots,\psi_{M}$. By combining (3.39) with (3.32), (3.38) and use (3.36), we get
$MN_{\rm gcd,\sigma}(F({\bf u}),G({\bf u}),r)+\frac{1}{u}N_{W,\sigma}(0,r)$
$\leq_{\operatorname{exc}}\left(\frac{w}{u}M-M+\epsilon\right)T_{\Psi,\sigma}(r)+M^{\prime}mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r)$
$+M\left(\frac{w}{u}M(b+2)+1\right)(T_{F,\sigma}(r)+T_{G,\sigma}(r))$
$+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+{1\over u}c_{{1\over W},\sigma}(r)+Mc_{{1\over h},\sigma}(r)+O(1).$
Finally, we need to estimate $T_{\Psi,\sigma}(r)$ to complete the proof. Recall that $\psi_{i}=\phi(\mathbf{u})/h$ and $\phi_{i}$ is of the form $F{\bf x}^{\bf i}$, or $G{\bf x}^{\bf i}$, which is a combination of monomial of degree $\leq m$ with coefficients coming from $F$ or $G$. Therefore, by (3.30)
$\log\|\psi_{i}\|\leq m\sum_{j=1}^{n}\log^{+}|u_{j}|+\log\|F\|+\log\|G\|+\log\frac{1}{h}.$
Combining this with Lemma 9 and (3.34), we obtain
$T_{\Psi,\sigma}(r)\leq mn\max_{1\leq i\leq n}T^{M}_{u_{i},\sigma}(r)+2\left(T_{F,\sigma}(r)+T_{G,\sigma}(r)\right)+c_{\frac{1}{h},\sigma}(r).$
@@
Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given as above with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Let $u_{1},\dots,u_{n}$, with $n\geq 2$, be meromorphic functions on $\mathbb{C}$} such that they are units on $\mathcal{Y}$}. Let $K$ be a subfield of the field of meromorphic functions on $\mathbb{C}$}. Let $F$ and $G$ be coprime polynomials in $K[x_{1},\ldots,x_{n}]$. Assume that at least one of the coefficients in the expansions of $F$ and $G$ is equal to 1. Let $\epsilon>0$. Then, by shrinking $r_{j}$ small enough (depending on $\epsilon$, $u_{i},1\leq i\leq n$, and on $F$ and $G$), we can choose a positive integer $m$ and positive reals $c_{i}$, $0\leq i\leq 5$, all depending only on $\epsilon$, $n$, and the degree of $F$ and $G$ such that, either
(3.42) $T^{M}_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\sigma}(r)\leq_{\rm exc}c_{1}\max_{1\leq i\leq n}\{\log T_{u_{i}}(r)\}+c_{2}(T_{F,\sigma}(r)+T_{G,\sigma}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
@@
(3.43) $N_{\rm gcd,\sigma}(F(u_{1},\ldots,u_{n}),G(u_{1},\ldots,u_{n}),r)$ $\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{4}(T_{F,\sigma}(r)+T_{G,\sigma}(r))$
$+c_{5}(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
@@
We note that $c_{i},1\leq i\leq 5,$ will be determined during the course of the proof. Recall that $M:=M_{m}:=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$ and $M^{\prime}:=M^{\prime}_{m}:=\binom{m+n}{n}-M=O(m^{n-2})$. Based on the above estimates, we can choose a real $C_{1}\geq 1$ independent of $\epsilon$ and the $u_{i},1\leq i\leq n,$ such that $m=C_{1}\epsilon^{-1}\geq 2d$, and